OpenStudy (anonymous):

A rectangular table has a length of 2x + 6 and a width of 2x − 5. What is its perimeter?

OpenStudy (austinl):

\[P = 2l + 2w\]

OpenStudy (austinl):

I am guiding you to the answer... we don't just give answers on this site.

OpenStudy (austinl):

l = length = 2x + 6 w = width = 2x - 5 \[P = 2l + 2w\rightarrow 2(2x+6) + 2(2x-5)=?\]

OpenStudy (austinl):

what is 2(2x+6)?

OpenStudy (austinl):

No.

OpenStudy (austinl):

a(b+c) = ab + ac So therefore 2(2x+6) = 4x + 12

OpenStudy (austinl):

No, you have to remember the signs. 4x - 10

OpenStudy (austinl):

\[P = 4x + 12 + 4x - 10\] Can you simplify that for me?

OpenStudy (austinl):

You CANNOT combine numbers and variables like that.

OpenStudy (austinl):

4x + 4x

OpenStudy (austinl):

12 + (-10)

OpenStudy (austinl):

Individually.